# Generalized open-loop Nash equilibria in linear-quadratic difference games with coupled-affine inequality constraints

### v4

@@ 1 Introduction
We denote the sets of natural numbers, real numbers, $n$-dimensional Euclidean space, $n$-dimensional non-negative orthant, and $n\times m$ real matrices by $\mathbb{N}$, $\mathbb{R}$, $\mathbb{R}^{n}$, +$\mathbb{R}^{n}_{+}$, and $\mathbb{R}^{n\times m}$, respectively. The transposes of a vector $a$ and a matrix $A$ are denoted by $a^{\prime}$ and $A^{\prime}$. The quadratic form $\|a\|_{A}$ is defined as $a^{\prime}Aa$. For $A\in\mathbb{R}^{n\times n}$ and $a\in\mathbb{R}^{n}$, where $n=n_{1}+\cdots+n_{K}$, $[A]_{ij}$ denotes the $n_{i}\times n_{j}$ submatrix, and $[a]_{i}$ denotes the $n_{i}\times 1$ subvector. Column vectors $[v_{1}^{\prime},\cdots,v_{n}^{\prime}]^{\prime}$ are written as $\mathrm{col}\{v_{1},\cdots,v_{n}\}$ or $\mathrm{col}\{v_{k}\}_{k=1}^{n}$. The identity and zero matrices of appropriate dimensions are denoted by $\mathbf{I}$ and $\mathbf{0}$, respectively. The block diagonal matrix with diagonal elements $M_{1},\cdots,M_{K}$ is denoted by $\oplus_{k=1}^{K}M_{k}$. The Kronecker product is $\otimes$. Vectors $x,y\in\mathbb{R}^{n}$ are complementary if $x\geq 0$, $y\geq 0$, and $x^{\prime}y=0$, denoted by $0\leq x\perp y\geq 0$.
@@ 3 Generalized Open-loop Nash Equilibrium
If $\Omega\neq\emptyset$, then from (4) and (5), we have that for every $x_{0}\in\mathsf{X}_{0}$, $R(x_{0})\neq\emptyset$. Due to the affine inequality constraints (2), $R(x_{0})$ is an intersection of $K$ closed half-spaces and, as a result, is a closed and convex subset of $\mathbb{R}^{Km}$ (with the usual topology on $\mathbb{R}^{Km}$). So item (i) ensures that $R(x_{0})$ is a closed, convex and bounded, which is required for the existence of GLONE. Item (ii) is required to satisfy the constraint qualification conditions. Item (iii) is a technical assumption required for obtaining individual players’ controls.
@@ 3.2 Sufficient conditions
where $E_{k}^{i}\in\mathbb{R}^{n\times n}$, $e_{k}^{i}\in\mathbb{R}^{n}$, $f_{k}^{i}\in\mathbb{R}$ for $k\in\mathsf{K}$. The next result, aimed at transforming DGC to a static game, is given as follows.
Theorem 1.
Let $\{\bm{x}^{\star},\bm{\lambda},\bm{\mu}^{\star}\}^{1}$ be a solution of LCS1. Let the solutions $\{E_{k}^{i},~{}k\in\mathsf{K}\}$ of the following symmetric difference matrix Riccati equations
$E_{k}^{i}$ $=A_{k}^{\prime}E_{k+1}^{i}A_{k}+Q^{i}_{k}-A_{k}^{\prime}E_{k+1}^{i}B_{k}^{i}\big{(}Y_{k}^{i}\big{)}^{-1}{B_{k}^{i}}^{\prime}E_{k+1}^{i}A_{k},$
$E_{K}^{i}$ $=Q^{i}_{K},$ (14)
exist (thus, the matrices $Y_{k}^{i}:=R^{ii}_{k}+{B_{k}^{i}}^{\prime}E_{k+1}^{i}B_{k}^{i},~{}k\in\mathsf{K}_{l}$ are invertible). Then for any strategy $(\mathsf{u}^{i},\mathsf{u}^{-i})\in R(x_{0})$, the difference equations
@@
with boundary conditions $e_{K}^{i}=p_{K}^{i}$, $b_{K}^{i}=0$ and $f_{K}^{i}=0$ are solvable for all $k\in\mathsf{K}$, where $\eta_{k}^{i}:=\sum_{j\in-i}B^{j}_{k}u_{k}^{j}$ and $\alpha_{k}^{i}:=\sum_{j\in-i}[N^{i}_{k}]_{j}u_{k}^{j}$. Further, the objective function of each player $i\in\mathsf{N}$ is given by
@@
Taking the telescopic sum of $V_{k+1}^{i}-V_{k}^{i}$ for all $k\in\mathsf{K}_{l}$, we obtain the relation (16).
@@ 3.2.1 Related static game with coupled constraints
Using Theorem 1, from DGC we derived an auxiliary static game with coupled constraints (SGCC) with Player $i$’s objective function given by (16) and the feasible strategy space of the players by (5). Next, we have the following assumption.
Assumption 2.
The symmetric matrix Riccati equations (14) admit solutions $\{E_{k}^{i},~{}k\in\mathsf{K},~{}i\in\mathsf{N}\}$ and thus the matrices in the set $\{Y_{k}^{i}=R^{ii}_{k}+{B_{k}^{i}}^{\prime}E_{k+1}^{i}B_{k}^{i},~{}k\in\mathsf{K}_{l},~{}i\in\mathsf{N}\}$ are invertible.
Let $(\bar{\mathsf{u}}^{i},\bar{\mathsf{u}}^{-i})\in R(x_{0})$ be a generalized Nash equilibrium associated with SGCC. This implies for each Player $i$ ($i\in\mathsf{N}$), $\bar{\mathsf{u}}^{i}$ solves the following constrained optimization problem
@@
${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\big{(}R^{ii}_{k}+{B_{k}^{i}}^{\prime}E_{k+1}^{i}B_{k}^{i}\big{)}\bar{u}_{k}^{i}+{B_{k}^{i}}^{\prime}E_{k+1}^{i}(A_{k}x_{k}^{\star}+\sum_{j\in-i}B^{j}_{k}\bar{u}_{k}^{j})}\\ {\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}-{B_{k}^{i}}^{\prime}E_{k+1}^{i}x_{k+1}^{\star}+{B_{k}^{i}}^{\prime}\lambda_{k+1}^{i}-[N^{i}_{k}]_{i}^{\prime}\mu_{k}^{i\star}=0.}$
@@ 3.2.2 Main result
Theorem 5 requires only that the matrices $\{Y_{k}^{i},~{}k\in\mathsf{K}_{l},~{}i\in\mathsf{N}\}$ are positive definite, which can be verified a priori using just the problem data related to state dynamics and objective functions, without solving the game. No assumption of positive semi-definiteness is made on $\{Q_{k}^{i},~{}k\in\mathsf{K},~{}i\in\mathsf{N}\}$. In fact, under Assumption 1.(iii), positive semi-definiteness of $\{Q_{k}^{i},~{}k\in\mathsf{K},~{}i\in\mathsf{N}\}$ is sufficient for the positive definiteness of $\{Y_{k}^{i},~{}k\in\mathsf{K}_{l},~{}i\in\mathsf{N}\}$.
@@ 3.3 Reformulation of LCS1 as a linear complementarity problem
In this section, under additional assumptions, we reformulate LCS1 (10) as a large-scale linear complementarity problem (LCP) using an approach similar to [19]. Initially, the state and co-state variables of LCS1 (10) are eliminated, allowing us to express these conditions implicitly in terms of the multipliers ($\bm{\mu}^{\star}$). We now make the following assumption.
@@ 4 Numerical Illustration
The open-loop potential game approach of [6] is applicable to the above flow control game only when $s^{1l}=s^{2l}$ and $S^{1l}=S^{2l}$ for $l=1,2$, as only in this case the sufficient conditions in [6, Lemma 3 and Lemma 4] are satisfied. Nevertheless, these constraints restrict the range of strategic scenarios within the flow-control game. Consequently, the approach presented in [6] is unsuitable for computing a GOLNE in such cases.
@@
It is easily verified that the conditions required in Theorems 6 are satisfied for the chosen parameter values. We used the freely available PATH solver (available at https://pages.cs.wisc.edu/ferris/path.html) to solve the LCP (41). The flow-rate constraints (44), with the above parameter values are given by

### v5

@@ 1 Introduction
We denote the sets of natural numbers, real numbers, $n$-dimensional Euclidean space, $n$-dimensional non-negative orthant, and $n\times m$ real matrices by $\mathbb{N}$, $\mathbb{R}$, $\mathbb{R}^{n}$, +$\mathbb{R}^{n}_{+}$, and $\mathbb{R}^{n\times m}$, respectively. The transposes of a vector $a$ and a matrix $A$ are denoted by $a^{\prime}$ and $A^{\prime}$. The quadratic form $a^{\prime}Aa$ is denoted by $\|a\|_{A}$. For $A\in\mathbb{R}^{n\times n}$ and $a\in\mathbb{R}^{n}$, where $n=n_{1}+\cdots+n_{K}$, $[A]_{ij}$ denotes the $n_{i}\times n_{j}$ submatrix, and $[a]_{i}$ denotes the $n_{i}\times 1$ subvector. Column vectors $[v_{1}^{\prime},\cdots,v_{n}^{\prime}]^{\prime}$ are written as $\mathrm{col}\{v_{1},\cdots,v_{n}\}$ or $\mathrm{col}\{v_{k}\}_{k=1}^{n}$. The identity and zero matrices of appropriate dimensions are denoted by $\mathbf{I}$ and $\mathbf{0}$, respectively. The block diagonal matrix with diagonal elements $M_{1},\cdots,M_{K}$ is denoted by $\oplus_{k=1}^{K}M_{k}$. The Kronecker product is $\otimes$. Vectors $x,y\in\mathbb{R}^{n}$ are complementary if $x\geq 0$, $y\geq 0$, and $x^{\prime}y=0$, denoted by $0\leq x\perp y\geq 0$.
@@ 3 Generalized Open-loop Nash Equilibrium
If $\Omega\neq\emptyset$, then from (4) and (5), we have that for every $x_{0}\in\mathsf{X}_{0}$, $R(x_{0})\neq\emptyset$. Due to the affine inequality constraints (2), $R(x_{0})$ is an intersection of $K$ closed half-spaces and, as a result, is a closed and convex subset of $\mathbb{R}^{Km}$ (with the usual topology on $\mathbb{R}^{Km}$). So item (i) ensures that $R(x_{0})$ is a closed, convex and bounded, which is required for the existence of GOLNE. Item (ii) is required to satisfy the constraint qualification conditions. Item (iii) is a technical assumption required for obtaining individual players’ controls.
@@ 3.2 Sufficient conditions
where $E_{k}^{i}\in\mathbb{R}^{n\times n}$, $e_{k}^{i}\in\mathbb{R}^{n}$, $f_{k}^{i}\in\mathbb{R}$ for $k\in\mathsf{K}$. Next, we have the following assumption on the matrices $\{E_{k}^{i},~{}k\in\mathsf{K},~{}i\in\mathsf{N}\}$.
Assumption 2.
The backward symmetric matrix Riccati equation
$E_{k}^{i}=A_{k}^{\prime}E_{k+1}^{i}A_{k}+Q^{i}_{k}-A_{k}^{\prime}E_{k+1}^{i}B_{k}^{i}\big{(}Y_{k}^{i}\big{)}^{-1}{B_{k}^{i}}^{\prime}E_{k+1}^{i}A_{k},~{}E_{K}^{i}=Q^{i}_{K},$ (14)
admits a solution $\{E_{k}^{i},~{}k\in\mathsf{K}\}$ for each $i\in\mathsf{N}$, and thus the matrices $\{Y_{k}^{i}=R^{ii}_{k}+{B_{k}^{i}}^{\prime}E_{k+1}^{i}B_{k}^{i},~{}k\in\mathsf{K}_{l},~{}i\in\mathsf{N}\}$ are invertible.
We note that if Assumption 2 holds and $\{\bm{x}^{\star},\bm{\lambda},\bm{\mu}^{\star}\}^{1}$ is a solution of LCS1, then for any strategy $(\mathsf{u}^{i},\mathsf{u}^{-i})\in R(x_{0})$, the difference equations
@@
with boundary conditions $e_{K}^{i}=p_{K}^{i}$, $b_{K}^{i}=0$, and $f_{K}^{i}=0$, are solvable for all $k\in\mathsf{K}$, where $\eta_{k}^{i}:=\sum_{j\in-i}B^{j}_{k}u_{k}^{j}$ and $\alpha_{k}^{i}:=\sum_{j\in-i}[N^{i}_{k}]_{j}u_{k}^{j}$. Using (14)–(15), the following auxiliary result expresses the objective function (1c) in a form that will be useful in deriving the sufficient condition later in Theorem 4.
Theorem 1.
Let Assumption 2 hold. Let $\{\bm{x}^{\star},\bm{\lambda},\bm{\mu}^{\star}\}^{1}$ be a solution of LCS1. Then, the objective function (1c) of each player $i\in\mathsf{N}$ can be expressed as:
@@
Taking the telescopic sum of $V_{k+1}^{i}-V_{k}^{i}$ for all $k\in\mathsf{K}_{l}$, we obtain the relation (16). ∎
@@ 3.2.1 Related static game with coupled constraints
Since the information structure is open-loop, Theorem 1 allows us to derive from DGC an auxiliary static game with coupled constraints (SGCC), in which Player $i$’s objective function is given by (16) and the players’ feasible strategy space is defined by (5); see also [26, 22] for the unconstrained case. Let $(\bar{\mathsf{u}}^{i},\bar{\mathsf{u}}^{-i})\in R(x_{0})$ be a generalized Nash equilibrium associated with SGCC. This implies for each Player $i$ ($i\in\mathsf{N}$), $\bar{\mathsf{u}}^{i}$ solves the following constrained optimization problem
@@
$\big{(}R^{ii}_{k}+{B_{k}^{i}}^{\prime}E_{k+1}^{i}B_{k}^{i}\big{)}\bar{u}_{k}^{i}+{B_{k}^{i}}^{\prime}E_{k+1}^{i}(A_{k}x_{k}^{\star}+\sum_{j\in-i}B^{j}_{k}\bar{u}_{k}^{j})\\ -{B_{k}^{i}}^{\prime}E_{k+1}^{i}x_{k+1}^{\star}+{B_{k}^{i}}^{\prime}\lambda_{k+1}^{i}-[N^{i}_{k}]_{i}^{\prime}\mu_{k}^{i\star}=0.$
@@ 3.2.2 Main result
Theorem 5 requires only that the matrices $\{Y_{k}^{i},~{}k\in\mathsf{K}_{l},~{}i\in\mathsf{N}\}$ be positive definite, which can be verified a priori using problem data given by (1), without solving the game. No assumption of positive semi-definiteness is made on $\{Q_{k}^{i},~{}k\in\mathsf{K},~{}i\in\mathsf{N}\}$. Under Assumption 1.(iii), the semi-definiteness of $\{Q_{k}^{i},~{}k\in\mathsf{K},~{}i\in\mathsf{N}\}$ is sufficient for the positive definiteness of $\{Y_{k}^{i},~{}k\in\mathsf{K}_{l},~{}i\in\mathsf{N}\}$.
@@ 3.3 Reformulation of LCS1 as a linear complementarity problem
In this section, under additional assumptions, we reformulate LCS1 (10) as a large-scale linear complementarity problem (LCP) using an approach similar to [19]. This approach eliminates the state and co-state variables of LCS1, allowing (10) to be expressed explicitly in terms of the multipliers ($\bm{\mu}^{\star}$). We now state the following assumption.
@@ 4 Numerical Illustration
The open-loop potential game approach of [6] is applicable to the above flow control game only when $s^{1l}=s^{2l}$ and $S^{1l}=S^{2l}$ for $l=1,2$, as only in this case are the sufficient conditions in [6, Lemma 3 and Lemma 4] satisfied. Nevertheless, these constraints restrict the range of strategic scenarios within the flow-control game. Consequently, the approach presented in [6] is unsuitable for computing a GOLNE in such cases.
@@
It is easily verified that the conditions required in Theorems 6 are satisfied for the chosen parameter values. We used the freely available PATH solver (available at https://pages.cs.wisc.edu/~ferris/path.html) to solve the LCP (41). The flow-rate constraints (44), with the above parameter values are given by